Freight car loadings over an 18-week period at a b
Answer #1

Linear regression

\hat{y} = a + bt

\hat{y} =  computed value of the variable to be predicted

a = y-axis intercept

b=slope of the regression line

t = independent variable

b= ( \sum ty - n\bar{t}\bar{y})/(\sum t^{2} - n\bar{t^{2}})

a= \bar{y}-b\bar{t}

n= no of months

From the above

\sum ty = 68445

\sum y   = 6465

\bar{t}  = 9.5

\bar{y} = 359.2

\sum t^{2}  = 2109

b = (68445-(18*359.2*9.5)/(2109-(18*9.5*9.5)

= 7027.5/484.5 =14.50

Slope b = 14.50

a = 359.2 - (14.5*9.5) = 221.4

\hat{y}=221.4+14.5 t

b) Forecast for week 20 = 221.4 + 14.5 *20 = 511.4

Forecast for week 21 = 221.4 + 14.5 *21 = 525.9

c)

Given Y= 840

then t = (840-221.4)/14.5 = 42.66

> This is 100% correct

Alex Richards Tue, Jan 24, 2023 7:59 PM